# SPD: Sync-Point Drop for efficient tensor parallelism of Large Language Models

### v1

@@ 4.2.1 Block-wise sync sensitivity identification
To identify the sensitivity of a block to SPD, we utilize the perplexity metric by measuring the relative impact of a block to performance (the difference from application between TP block and SPD block in Figure 4) as sensitivity measurement. For example, when we measure the sensitivity of $i$-th block to SPD, we apply SPD to all blocks starting from the {$i+1$}-th block to the final block and measure the perplexity, while leaving the $i$-th block unchanged (TP block). Then we measure the perplexity by additionally modifying the system setting of $i$-th block to SPD. The difference in perplexity before and after applying SPD to $i$-th block is used as a measure of sensitivity. In this measurement, we use calibration data obtained by sampling a small portion of the large training dataset. By progressive replacement of TP block to SPD block and measurement of quality degradation as relative perplexity difference, we can compare the sensitivity between blocks in the entire model and classify the blocks into three sensitivity categories (ISB, SB, and ESB).
@@ 5 Experiments
Figure 7: LLaMA2 distributed inference average accuracy on zero-shot tasks (top) and normalized latency speedup (bottom). Speedup is normalized based on latency of 0% (No SPD state which consists of TP blocks in entire model) in each distributed inference setting. ‘ZS’ represents applying zero-shot dropping to all blocks. ‘ZS+B2B’ represents applying zero-shot dropping on ISBs and block-to-block distillation to the other remaining SBs and ESBs. ‘ZS+B2B+HG’ is applying zero-shot dropping on ISBs and block-to-block distillation to SBs and block-to-block distillation with head grouping initialization to the other remaining ESBs. ‘HBW’ represents high GPU interconnect bandwidth of 300GB/s setting and ‘LBW’ represents low GPU interconnect bandwidth of 10GB/s setting.

### v3

@@ 4.2.1 Block-wise sync sensitivity identification
To identify the sensitivity of a block to SPD, we utilize the perplexity metric by measuring the relative impact of a block to performance (the difference from application between TP block and SPD block in Figure 4) as sensitivity measurement. For example, when we measure the sensitivity of $i$-th block to SPD, we apply SPD to all blocks starting from the {$i+1$}-th block to the final block and measure the perplexity, while leaving the $i$-th block unchanged (TP block). We then measure the perplexity by additionally modifying the system setting of $i$-th block to SPD. The difference in perplexity before and after applying SPD to $i$-th block is used as a measure of sensitivity. In this measurement, we use calibration data obtained by sampling a small portion of the large training dataset. By progressive replacement of TP block to SPD block and measurement of quality degradation as relative perplexity difference, we can compare the sensitivity between blocks in the entire model and classify the blocks into three sensitivity categories (ISB, SB, and ESB).
@@ 5 Experiments
Figure 7: LLaMA2 distributed inference average accuracy on zero-shot tasks (top) and normalized speedup of time to first token (bottom). Speedup is normalized based on latency of 0% (No SPD state which consists of TP blocks in entire model) in each distributed inference setting. ‘ZS’ represents applying zero-shot dropping to all blocks. ‘ZS+B2B’ represents applying zero-shot dropping on ISBs and block-to-block distillation to the other remaining SBs and ESBs. ‘ZS+B2B+HG’ is applying zero-shot dropping on ISBs and block-to-block distillation to SBs and block-to-block distillation with head grouping initialization to the other remaining ESBs. ‘HBW’ represents high GPU interconnect bandwidth of 300GB/s setting and ‘LBW’ represents low GPU interconnect bandwidth of 10GB/s setting.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgments
We thank Rasoul Shafipour, Raviteja Vemulapalli, Chenfan Sun, Iman Mirzadeh, Keivan Alizadeh Vahid, Devang Naik, Kulin Seth, and Brendan Langoulant for valuable discussions and comments.
@@ B.1 Effects of design choice in block design
Appendix C Discussion
C.1 Sensitivity identification
In Section 4.2.1, when measuring the sensitivity of a block to SPD, we apply SPD to consecutive blocks starting from the selected block and extending toward the last block. This approach allows us to evaluate the worst-case impact of SPD on the block’s output while ensuring that its input activations remain numerically identical to the original, unaffected by prior SPD modifications. By analyzing the perplexity difference between a given block and the next block, we can isolate the effect of SPD at that specific layer, excluding any influence from subsequent layers. This method provides a fast yet precise and independent measurement of each block’s sensitivity.
C.2 Compatibility with other parallelism
SPD can be integrated with diverse settings to suit specific application contexts. In this section, we describe how SPD can be combined with such settings.
Data parallelism Data parallelism can be achieved with SPD by replicating an SPD model across distributed GPU environments. For example, in an 8-GPUs configuration, a 4-GPUs SPD model can be instantiated as a single replica, which is then duplicated across the remaining 4-GPUs. Consequently, two replicas operate in parallel, forming a data parallelism setup with SPD.
Pipeline parallelism Pipeline parallelism can be incorporated into SPD by applying it at the level of each parallelized decoder block, in which the attention and feed-forward network (FFN) components are already distributed across multiple GPUs. For instance, in an SPD configuration where a single decoder block is partitioned across 4-GPUs, pipeline parallelism can be realized by further dividing the entire distributed blocks into two sequential groups. This results in a 4×2-GPUs system, where each of the four SPD branches is executed in a two-stage pipeline across a pair of GPUs, thereby enabling pipeline parallelism within SPD blocks.
Hybrid parallelism A pipeline-parallelized model can be instantiated as a replica to enable data parallelism. By replicating this pipeline-parallelized model, data parallelism can be applied on top of the pipeline setup. This approach ultimately facilitates the implementation of combined parallelism in conjunction with SPD.
C.3 Impact of SPD on tensor parallelism compared to pipeline parallelism
In this work, we focus on optimizing tensor parallelism (TP) by addressing the communication bottlenecks that frequently arise in distributed inference systems. Although TP is widely adopted in practice due to its engineering simplicity and high compute utilization, pipeline parallelism (PP) is often considered a promising alternative, primarily due to its lower communication overhead. Nonetheless, TP remains the preferred strategy in many real-world distributed inference deployments, even under constrained interconnect bandwidth, due to its ability to more effectively utilize available compute resources.
While the relative reduction in communication volume achieved by SPD may appear modest (maximum 50% since SPD only targets the sync-point after self attention output), its impact on end-to-end latency is substantial. This is because the reduction in communication bottlenecks allows the remaining computational tasks, which are less affected by the system’s bandwidth, to proceed more efficiently, ultimately improving the end-to-end performance. Therefore, TP optimized by SPD is more effective in deployment scenarios where compute resources need to be fully utilized across GPUs, as seen in prefill stage of large language model (Figure 7).